OpenStudy (anonymous):

What is the common ratio between successive terms in the sequence? 27, 9, 3, 1,1/3 , 1/9, 1/27, ... A. –3 B. –1/3 C. 1/3 D. 3

OpenStudy (phi):

the are asking what do you multiply 27 by to get 9? 27 * x = 9 where x is the unknown number if you divide both sides by 27 you get x= 9/27 you can simplify that fraction to get the answer.

OpenStudy (phi):

or what do you multiply 9 by to get 3? or what do you multiply 3 by to get 1 ?

OpenStudy (anonymous):

3 right? @phi

OpenStudy (phi):

are you saying 27*3 = 9 ?

OpenStudy (danjs):

common ratio of a geometric series r, the ratio of (n+1)/n term

OpenStudy (danjs):

every term divided by the previous term should give you the same number as the common ratio

OpenStudy (danjs):

9 / 27 3 / 9 1/3 .....

OpenStudy (danjs):

all successive have common ratio r = 1/3

OpenStudy (anonymous):

thank you @DanJS

jaynator495 (jaynator495):

Welcome To OpenStudy! Here you will find great helpers and friends, a community of students that help students! We hope you enjoy the experience!